OpenStudy (ccastudentgrade10):

write an explicit formula for the sequence -4, 1, 6, 11

OpenStudy (ccastudentgrade10):

a. an=-4+5(n+1) b. an=-4+5(n=1) c.an=-4+(n-1) d.an=-4+5n

OpenStudy (mathmale):

You'll need to find the "common difference." What number, when added to an element of this sequence, correctly produces the next number in the sequence?